# AGZO: Activation-Guided Zeroth-Order Optimization for LLM Fine-Tuning

### v1

@@ 6.1 Alignment to the true gradient
This diagnostic experiment is designed to validate the directional-quality analysis in Section 5. In particular, Section 5 derives a closed-form expression for the expected cosine similarity of AGZO in the noiseless regime (Theorem 5.4) and shows that, under activation spectral concentration, AGZO achieves a strictly larger expected cosine similarity than MEZO (Theorem 5.7).
To empirically test this prediction in a realistic fine-tuning setting, we fine-tune Qwen3-0.6B on SST-2 and track, at training step $t$, the cosine similarity between the ZO estimated gradient and the exact backpropagation gradient computed on the same mini-batch. As shown in Figure 2, AGZO consistently yields higher cosine similarity than MEZO throughout training. This observation directly supports the theoretical conclusion that activation-guided low-rank perturbations improve directional alignment relative to dense Gaussian ZO perturbations.
6.2 End-to-end fine-tuning performance
Qwen3-0.6B.
Table 1 shows that AGZO achieves consistently stronger downstream performance than existing ZO baselines across a broad set of benchmarks. This improvement is aligned with the mechanism validated in Section 6.1: by producing update directions that are better aligned with the true gradient, AGZO enables more effective optimization under the same query budget. As a result, AGZO converges to better solutions and noticeably narrows the performance gap between zeroth-order fine-tuning and first-order training.
Pangu-1B.
Table 2 summarizes the performance of AGZO and various baselines on the openPangu-embedded-1B model across multiple downstream tasks. Overall, AGZO consistently outperforms existing zeroth-order baselines (MEZO and LOZO) and non-training baselines (zero-shot prompting and in-context learning) on most tasks, demonstrating the effectiveness of our approach in adapting large models with limited gradient information. The average performance shows that AGZO achieves 0.720, surpassing the best ZO baseline (MEZO) by 1.4%.
@@ 6.3 Peak GPU memory footprint
We empirically validate this memory analysis by measuring the peak GPU memory footprint when fine-tuning Qwen3-0.6B and Pangu-1B DROP task. We sweep the two primary drivers of training-time memory: sequence length and batch size. Figure 3 reports two representative slices. In Figures 3(a) and 4(a), we fix the batch size to 4 and increase the sequence length. FO (backpropagation) exhibits rapidly growing memory usage and becomes out-of-memory (OOM) at long contexts, whereas AGZO/MEZO/LOZO remain substantially lower and continue to run. In Figures 3(b) and 4(b), we fix the sequence length to 256 and increase the batch size. FO again hits OOM at moderate batch sizes, while ZO methods remain feasible for significantly larger batches.
@@ 7 Conclusions
We introduced AGZO, a zeroth-order fine-tuning method that leverages per-iteration activation structure to construct low-rank, activation-guided perturbations for linear layers. By extracting compact activation subspaces on the fly via lightweight power iteration, AGZO concentrates ZO updates on directions that are more tightly coupled to backpropagation signals without storing activations across iterations. We provided a principled explanation by interpreting AGZO as optimizing a subspace-smoothed objective and showing that, under activation spectral concentration, its update direction achieves a strictly larger expected cosine similarity to the true gradient than prior ZO baselines. Experiments on Qwen3 models under practical GPU memory constraints corroborate the analysis, demonstrating improved downstream performance and directional fidelity over existing zeroth-order methods while preserving the favorable memory profile of ZO fine-tuning.
@@ Acknowledgements
This work is supported in part by funding from Huawei (TC20250928057) and CUHK (4937007, 4937008, 5501329, 5501517).
@@ A.1.1 Smoothing operators
Fix a step radius $\mu>0$. For a given layer $l$, let $A_{l}\in\mathbb{R}^{n_{l}\times r_{l}}$ have orthonormal columns ($A_{l}^{\top}A_{l}=I_{r_{l}}$), and let $R_{l}\in\mathbb{R}^{m_{l}\times r_{l}}$ have i.i.d. $\mathcal{N}(0,1)$ entries. Define the rank-$r_{l}$ matrix direction $\Delta_{l}=R_{l}A_{l}^{\top}$ and the block direction $\Delta=\{\Delta_{l}\}_{l=1}^{L}$.
@@
For MEZO, let $U_{l}\in\mathbb{R}^{m_{l}\times n_{l}}$ have i.i.d. $\mathcal{N}(0,1)$ entries and set $\Delta_{l}=U_{l}$ (full-dimensional). Define
@@ A.1.2 Subspace-smoothing gradient identity for AGZO
Let $R\in\mathbb{R}^{m\times r}$ have i.i.d. $\mathcal{N}(0,1)$ entries and let $M\in\mathbb{R}^{m\times r}$ be deterministic. Then
@@
Similarly, if $U\in\mathbb{R}^{m\times n}$ is i.i.d. standard Gaussian and $G\in\mathbb{R}^{m\times n}$, then $\mathbb{E}\,[\langle G,U\rangle\,U]=G$.
@@
We proceed entrywise. For $a\in\{1,\dots,m\}$ and $b\in\{1,\dots,r\}$,
$\big[\mathbb{E}\,\big[\langle M,R\rangle\,R\big]\big]_{ab}=\mathbb{E}\Big[\Big(\sum_{i=1}^{m}\sum_{j=1}^{r}M_{ij}R_{ij}\Big)R_{ab}\Big]=\sum_{i=1}^{m}\sum_{j=1}^{r}M_{ij}\,\mathbb{E}\big[R_{ij}R_{ab}\big].$
@@
The isotropic version is identical: for $U\in\mathbb{R}^{m\times n}$ i.i.d. $\mathcal{N}(0,1)$ and deterministic $G$,
$\big[\mathbb{E}\,[\langle G,U\rangle\,U]\big]_{ab}=\sum_{i=1}^{m}\sum_{j=1}^{n}G_{ij}\,\mathbb{E}[U_{ij}U_{ab}]=G_{ab},$
@@
Vectorized view. Let $z=\mathrm{vec}(R)\sim\mathcal{N}(0,I_{mr})$ and $x=\mathrm{vec}(M)$. Then $\mathbb{E}[(x^{\top}z)z]=\mathbb{E}[zz^{\top}]x=I_{mr}x=x$, which is the same identity reshaped to matrices. ∎
@@
Let $R\in\mathbb{R}^{m\times r}$ have i.i.d. $\mathcal{N}(0,1)$ entries with joint density $p(R)=\prod_{a=1}^{m}\prod_{b=1}^{r}\phi(R_{ab})$, where $\phi(z)=(2\pi)^{-1/2}e^{-z^{2}/2}$. Let $h:\mathbb{R}^{mr}\to\mathbb{R}$ be $C^{1}$ with $\mathbb{E}\,|h(R)|<\infty$ and $\mathbb{E}\,\|\nabla_{R}h(R)\|_{F}<\infty$. Assume moreover the following sub-Gaussian growth condition:
@@
Fix $W$ and a batch $B$. Let $A_{l}\in\mathbb{R}^{n_{l}\times r_{l}}$ have orthonormal columns, and let $R_{l}\in\mathbb{R}^{m_{l}\times r_{l}}$ have i.i.d. $\mathcal{N}(0,1)$ entries, independently across $l$. Define $\Delta_{l}:=R_{l}A_{l}^{\top}$, $\Delta:=\{\Delta_{l}\}_{l=1}^{L}$,
@@
By Lemma A.2,
@@ A.1.3 Isotropic Gaussian smoothing identity for MEZO
] Fix $W$ and a batch $B$. For each layer $l$, let $U_{l}\in\mathbb{R}^{m_{l}\times n_{l}}$ have i.i.d. $\mathcal{N}(0,1)$ entries, independently across $l$, and define the full-direction block $\Delta_{l}:=U_{l}$ and $\Delta:=\{\Delta_{l}\}_{l=1}^{L}$. Define
@@ A.1.4 Consequences and specializations
Let $S_{l}^{(r_{l})}=\mathrm{span}(U_{l}^{(r_{l})})$ be the leading activation subspace and suppose $A_{l}$ is an orthonormal basis that (approximately) spans $S_{l}^{(r_{l})}$. By Lemma A.3,
@@
In the ideal alignment case $S_{l}^{(r_{l})}=\mathrm{col}(H_{l})$, using $\mathrm{row}(\nabla_{W_{l}}F)\subseteq\mathrm{col}(H_{l})$ we have $\nabla_{W_{l}}F(W)=\nabla_{W_{l}}F(W)\,\Pi_{S_{l}^{(r_{l})}}$, so the only bias comes from smoothing.
@@ A.2.1 Expected Cosine Similarity for AGZO
For simplification, we denote the true gradient as $G\in\mathbb{R}^{m\times n}$ and the agzo approximated gradient as $\widehat{G}$. Let $A\in\mathbb{R}^{n\times r}$ have orthonormal columns ($A^{\top}A=I_{r}$) and let $R\in\mathbb{R}^{m\times r}$ have i.i.d. $\mathcal{N}(0,1)$ entries. In AGZO we perturb with
@@
$\mathbb{E}_{R}\!\left[\cos\!\big(\widehat{G}_{0},\,G\big)\right],\qquad\cos(X,G):=\frac{\langle X,G\rangle}{\|X\|_{F}\,\|G\|_{F}}.$
@@
Let $G\in\mathbb{R}^{m\times n}$ be fixed and $A\in\mathbb{R}^{n\times r}$ have orthonormal columns. Let $\widehat{G}^{\mathrm{AGZO}}_{0}$ be the noiseless AGZO estimator constructed from $\Delta=RA^{\top}$ with $R\sim\mathcal{N}(0,I_{m\times r})$. Then
$\mathbb{E}_{R}\big[\cos\big(\widehat{G}^{\mathrm{AGZO}}_{0},G\big)\big]=\beta_{mr}\,\frac{\|GA\|_{F}}{\|G\|_{F}},$ (48)
@@
depends only on the product dimension $mr$. Equivalently,
@@
and $\beta_{D}$ decays on the order of $1/\sqrt{D}$ as $D$ increases.
@@
Step 4 (vectorization and rotational reduction). Let $r:=\mathrm{vec}(R)\in\mathbb{R}^{d}$ with $d=mr$ and note $r\sim\mathcal{N}(0,I_{d})$; also set $k:=\mathrm{vec}(GA)$, so $\langle R,GA\rangle=r^{\top}k$ and $\|R\|_{F}=\|r\|_{2}$. Thus
@@
$\mathbb{E}_{R}\!\left[\frac{|r^{\top}\hat{k}|}{\|r\|_{2}}\right]=\mathbb{E}\big[\,|U_{1}|\,\big]=:\beta_{mr},$
where $U=(U_{1},\dots,U_{mr})$ is uniform on $\mathbb{S}^{mr-1}$. Therefore
$\mathbb{E}_{R}\!\left[\cos\!\big(\widehat{G}_{0},\,G\big)\right]=\frac{\|k\|_{2}}{\|G\|_{F}}\,\beta_{mr}=\beta_{mr}\,\frac{\|GA\|_{F}}{\|G\|_{F}}.$
Step 5 (closed form for $\beta_{mr}$). The marginal density of $U_{1}$ is
$f_{mr}(t)=c_{mr}\,(1-t^{2})^{\frac{mr-3}{2}},\quad t\in[-1,1],\qquad c_{mr}=\frac{\Gamma(\tfrac{mr}{2})}{\sqrt{\pi}\,\Gamma(\tfrac{mr-1}{2})}.$
@@
$\beta_{mr}=\mathbb{E}|U_{1}|=2\!\int_{0}^{1}t\,f_{mr}(t)\,dt=2c_{mr}\!\int_{0}^{1}t\,(1-t^{2})^{\frac{mr-3}{2}}\,dt.$
@@
$\beta_{mr}=c_{mr}\!\int_{0}^{1}(1-u)^{\frac{mr-3}{2}}\,du=c_{mr}\cdot\frac{2}{mr-1}=\frac{\Gamma(\tfrac{mr}{2})}{\sqrt{\pi}\,\Gamma(\tfrac{mr+1}{2})}.$
@@
In particular, $\beta_{mr}\approx c/\sqrt{mr}$ and $\beta_{mn}\approx c/\sqrt{mn}$ for a constant $c\approx\sqrt{2/\pi}$, so working in a much smaller subspace (with $r\ll n$) already improves the expected cosine similarity by a factor around $\sqrt{n/r}$.
@@
By Gautschi’s inequality with $s=\tfrac{1}{2}$ applied at $u-\tfrac{1}{2}$ (so $u>\tfrac{1}{2}$):
@@
Wendel’s inequality (1948) states that for $u>0$ and $s\in(0,1)$,
$\frac{\Gamma(u+s)}{u^{s}\,\Gamma(x)}\;\leq\;1.$

### v2

@@ 6.1 Alignment to the True Gradient
This diagnostic experiment is designed to validate the gradient accuracy analysis in Section 5, which shows that under activation spectral concentration, AGZO achieves a strictly larger expected cosine similarity than MEZO.
To empirically test this prediction in a realistic fine-tuning setting, we fine-tune Qwen3-0.6B with Testbed One on SST-2 and track, at training step $t$, the cosine similarity between the ZO estimated gradient and the exact backpropagation gradient computed on the same mini-batch. As shown in Figure 2, AGZO consistently yields higher cosine similarity than MEZO throughout training. This observation directly supports the theoretical conclusion that activation-guided low-rank perturbations improve directional alignment relative to dense Gaussian ZO perturbations.
6.2 End-to-End Fine-Tuning Performance (Testbed One)
Qwen3-0.6B. Table 1 shows that AGZO achieves consistently stronger downstream performance than existing ZO baselines across a broad set of benchmarks. By producing update directions that are better aligned with the true gradient, AGZO enables more effective optimization under the same query budget. As a result, AGZO converges to better solutions and noticeably narrows the performance gap between zeroth-order fine-tuning and first-order training.
Qwen3-4B. Table 2 reports the results on Qwen3-4B. Under the same hardware setting, FO method runs out of memory, whereas ZO methods remain feasible. AGZO consistently outperforms MeZO and LOZO on this larger scale. AGZO successfully bridges the gap between memory efficiency and optimization quality, enabling effective fine-tuning of larger models on consumer-grade GPUs.
Pangu-1B. Table 3 summarizes the performance of AGZO and various baselines on the Pangu-1B model. Overall, AGZO consistently outperforms existing zeroth-order baselines (MEZO and LOZO) and non-training baselines (zero-shot prompting and in-context learning) on most tasks, demonstrating the effectiveness of our approach in adapting large models with limited gradient information.
6.3 End-to-End Cross-Platform Fine-Tuning Performance
We evaluate the cross-platform inference performance, i.e, evaluating the performance on NPU (with Testbed Two) with GPU-trained models (with Testbed One). Table 4 presents the performance of AGZO and various baselines on openPangu-embedded-1B across on NPU. Across both GPU and NPU, AGZO consistently achieves the best results among zeroth-order methods and non-training baselines on most downstream tasks. On the NPU, AGZO attains an average score of 0.709, outperforming other ZO baselines on tasks such as SST2, COPA, BoolQ, RTE, and WiC. The slightly lower performance on the NPU compared to the GPU may be attributed to subtle differences in numerical precision, memory layout, or low-level kernel implementations, which can affect the propagation of small perturbations used in zeroth-order optimization.
Task FO AGZO MEZO LOZO Zero ICL
SST2 0.821 0.765 0.766 0.718 0.571 0.710
COPA 0.800 0.770 0.740 0.720 0.760 0.740
CB 0.696 0.696 0.732 0.643 0.482 0.446
BoolQ 0.752 0.728 0.697 0.694 0.696 0.731
RTE 0.780 0.729 0.729 0.682 0.578 0.682
WiC 0.657 0.567 0.552 0.542 0.469 0.495
Avg. 0.738 0.709 0.703 0.667 0.593 0.636
@@ 6.4 Peak GPU Memory Footprint
We empirically validate this memory analysis by measuring the peak GPU memory footprint when fine-tuning Qwen3-0.6B on DROP task with Testbed one. We sweep the two primary drivers of training-time memory: sequence length and batch size. In Figure 3(a), we fix the batch size to 4 and increase the sequence length. FO exhibits rapidly growing memory usage and becomes out-of-memory (OOM) at long contexts, whereas ZO methods remain substantially lower and continue to run. In Figure 3(b), we fix the sequence length to 256 and increase the batch size. FO again hits OOM at moderate batch sizes, while ZO methods remain feasible for significantly larger batches.
@@ 7 Conclusions
In this paper, we propose AGZO, a zeroth-order fine-tuning method that leverages per-iteration activation structure to construct low-rank, activation-guided perturbations for linear layers. By extracting compact activation subspaces on the fly via lightweight power iteration, AGZO concentrates ZO updates on directions that are intrinsically coupled to backpropagation signals, all without storing activations across iterations. Theoretically, we prove that under activation spectral concentration, the AGZO update direction achieves a strictly larger expected cosine similarity to the true gradient than prior isotropic ZO baselines. Extensive experiments on Qwen3 and Pangu models corroborates this analysis, demonstrating that AGZO consistently outperforms existing zeroth-order methods in both downstream performance and directional fidelity, while maintaining the strict memory efficiency required for fine-tuning large language models.
@@ Acknowledgements
This work is supported in part by funding from Huawei (TC20250928057) and CUHK (4937007, 4937008, 5501329, 5501517). We thank Yu Pan for useful discussions.
@@ A.1.1 Smoothing operators
Fix a step radius $\mu>0$. For a given layer $l$, let $A_{l}\in\mathbb{R}^{d_{in}\times r}$ have orthonormal columns ($A_{l}^{\top}A_{l}=I_{r}$), and let $R_{l}\in\mathbb{R}^{d_{out}\times r}$ have i.i.d. $\mathcal{N}(0,1)$ entries. Define the rank-$r$ matrix direction $\Delta_{l}=R_{l}A_{l}^{\top}$ and the block direction $\Delta=\{\Delta_{l}\}_{l=1}^{L}$.
@@
For MEZO, let $U_{l}\in\mathbb{R}^{d_{out}\times d_{in}}$ have i.i.d. $\mathcal{N}(0,1)$ entries and set $\Delta_{l}=U_{l}$ (full-dimensional). Define
@@ A.1.2 Subspace-smoothing gradient identity for AGZO
Let $R\in\mathbb{R}^{d_{out}\times r}$ have i.i.d. $\mathcal{N}(0,1)$ entries and let $M\in\mathbb{R}^{d_{out}\times r}$ be deterministic. Then
@@
Similarly, if $U\in\mathbb{R}^{d_{out}\times d_{in}}$ is i.i.d. standard Gaussian and $G\in\mathbb{R}^{d_{out}\times d_{in}}$, then $\mathbb{E}\,[\langle G,U\rangle\,U]=G$.
@@
We proceed entrywise. For $a\in\{1,\dots,d_{out}\}$ and $b\in\{1,\dots,r\}$,
$\big[\mathbb{E}\,\big[\langle M,R\rangle\,R\big]\big]_{ab}=\mathbb{E}\Big[\Big(\sum_{i=1}^{d_{out}}\sum_{j=1}^{r}M_{ij}R_{ij}\Big)R_{ab}\Big]=\sum_{i=1}^{d_{out}}\sum_{j=1}^{r}M_{ij}\,\mathbb{E}\big[R_{ij}R_{ab}\big].$
@@
The isotropic version is identical: for $U\in\mathbb{R}^{d_{out}\times d_{in}}$ i.i.d. $\mathcal{N}(0,1)$ and deterministic $G$,
$\big[\mathbb{E}\,[\langle G,U\rangle\,U]\big]_{ab}=\sum_{i=1}^{d_{out}}\sum_{j=1}^{d_{in}}G_{ij}\,\mathbb{E}[U_{ij}U_{ab}]=G_{ab},$
@@
Vectorized view. Let $z=\mathrm{vec}(R)\sim\mathcal{N}(0,I_{d_{out}r})$ and $x=\mathrm{vec}(M)$. Then $\mathbb{E}[(x^{\top}z)z]=\mathbb{E}[zz^{\top}]x=Ix=x$, which is the same identity reshaped to matrices. ∎
@@
Let $R\in\mathbb{R}^{d_{out}\times r}$ have i.i.d. $\mathcal{N}(0,1)$ entries with joint density $p(R)=\prod_{a=1}^{d_{out}}\prod_{b=1}^{r}\phi(R_{ab})$, where $\phi(z)=(2\pi)^{-1/2}e^{-z^{2}/2}$. Let $h:\mathbb{R}^{d_{out}r}\to\mathbb{R}$ be $C^{1}$ with $\mathbb{E}\,|h(R)|<\infty$ and $\mathbb{E}\,\|\nabla_{R}h(R)\|_{F}<\infty$. Assume moreover the following sub-Gaussian growth condition:
@@
Fix $W$ and a batch $B$. Let $A_{l}\in\mathbb{R}^{d_{in}\times r}$ have orthonormal columns, and let $R_{l}\in\mathbb{R}^{d_{out}\times r}$ have i.i.d. $\mathcal{N}(0,1)$ entries, independently across $l$. Define $\Delta_{l}:=R_{l}A_{l}^{\top}$, $\Delta:=\{\Delta_{l}\}_{l=1}^{L}$,
@@
Note that the L-Lipschitz gradient assumption implies at most quadratic growth of $f$, which satisfies Condition (A) required by Lemma A.2. Hence we have,
@@ A.1.3 Isotropic Gaussian smoothing identity for MEZO
] Fix $W$ and a batch $B$. For each layer $l$, let $U_{l}\in\mathbb{R}^{d_{out}\times d_{in}}$ have i.i.d. $\mathcal{N}(0,1)$ entries, independently across $l$, and define the full-direction block $\Delta_{l}:=U_{l}$ and $\Delta:=\{\Delta_{l}\}_{l=1}^{L}$. Define
@@ A.1.4 Consequences and specializations
Let $S_{l}^{(r)}=\mathrm{span}(U_{l}^{(r)})$ be the leading activation subspace and suppose $A_{l}$ is an orthonormal basis that (approximately) spans $S_{l}^{(r)}$. By Lemma A.3,
@@
In the ideal alignment case $S_{l}^{(r)}=\mathrm{col}(H_{l})$, using $\mathrm{row}(\nabla_{W_{l}}F)\subseteq\mathrm{col}(H_{l})$ we have $\nabla_{W_{l}}F(W)=\nabla_{W_{l}}F(W)\,\Pi_{S_{l}^{(r)}}$, so the only bias comes from smoothing.
@@ A.2.1 Expected Cosine Similarity for AGZO
For simplification, we denote the true gradient as $G\in\mathbb{R}^{d_{out}\times d_{in}}$ and the agzo approximated gradient as $\widehat{G}$. Let $A\in\mathbb{R}^{d_{in}\times r}$ have orthonormal columns ($A^{\top}A=I_{r}$) and let $R\in\mathbb{R}^{d_{out}\times r}$ have i.i.d. $\mathcal{N}(0,1)$ entries. In AGZO we perturb with
@@
The AGZO estimator for layer $\ell$ can be written as
where
$\phi(W,\Delta;B)=\frac{f(W+\mu\Delta,B)-f(W,B)}{\mu}.$ (37)
@@
$\mathbb{E}_{R}\!\left[\cos\!\big(\widehat{G}_{0},\,G\big)\right],\qquad\cos(\widehat{G}_{0},G):=\frac{\langle\widehat{G}_{0},G\rangle}{\|\widehat{G}_{0}\|_{F}\,\|G\|_{F}}.$
@@
Let $\widehat{G}^{\mathrm{AGZO}}_{0}$ be the noiseless AGZO estimator constructed from $\Delta=RA^{\top}$ with $R\sim\mathcal{N}(0,I_{d_{out}\times r})$. Then
$\mathbb{E}_{R}\big[\cos\big(\widehat{G}^{\mathrm{AGZO}}_{0},G\big)\big]=\beta_{d_{out}r}\,\frac{\|GA\|_{F}}{\|G\|_{F}},$ (38)
@@
depends only on the product dimension $d_{out}r$. Equivalently,
@@
and for any $D\geq 2$, $\beta_{D}$ satisfies the tight bounds:
$\sqrt{\frac{2}{\pi D}}\;\leq\;\beta_{D}\;\leq\;\sqrt{\frac{2}{\pi(D-1)}}.$ (41)
@@
Step 4 (vectorization and rotational reduction). Let $r:=\mathrm{vec}(R)\in\mathbb{R}^{d}$ with $d=d_{out}r$ and note $r\sim\mathcal{N}(0,I_{d})$; also set $k:=\mathrm{vec}(GA)$, so $\langle R,GA\rangle=r^{\top}k$ and $\|R\|_{F}=\|r\|_{2}$. Thus
@@
$\mathbb{E}_{R}\!\left[\frac{|r^{\top}\hat{k}|}{\|r\|_{2}}\right]=\mathbb{E}\big[\,|U_{1}|\,\big]=:\beta_{d_{out}r},$
where $U=(U_{1},\dots,U_{mr})$ is uniform on $\mathbb{S}^{d_{out}r-1}$. Therefore
$\mathbb{E}_{R}\!\left[\cos\!\big(\widehat{G}_{0},\,G\big)\right]=\frac{\|k\|_{2}}{\|G\|_{F}}\,\beta_{d_{out}r}=\beta_{d_{out}r}\,\frac{\|GA\|_{F}}{\|G\|_{F}}.$
Step 5 (closed form for $\beta_{d_{out}r}$). The marginal density of $U_{1}$ is
$f_{d_{out}r}(t)=c_{d_{out}r}\,(1-t^{2})^{\frac{d_{out}r-3}{2}},\quad t\in[-1,1],\qquad c_{d_{out}r}=\frac{\Gamma(\tfrac{d_{out}r}{2})}{\sqrt{\pi}\,\Gamma(\tfrac{d_{out}r-1}{2})}.$
@@
$\beta_{d_{out}r}=\mathbb{E}|U_{1}|=2\!\int_{0}^{1}t\,f_{d_{out}r}(t)\,dt=2c_{d_{out}r}\!\int_{0}^{1}t\,(1-t^{2})^{\frac{d_{out}r-3}{2}}\,dt.$
@@
$\beta_{d_{out}r}=c_{d_{out}r}\!\int_{0}^{1}(1-u)^{\frac{d_{out}r-3}{2}}\,du=c_{d_{out}r}\cdot\frac{2}{d_{out}r-1}=\frac{\Gamma(\tfrac{d_{out}r}{2})}{\sqrt{\pi}\,\Gamma(\tfrac{d_{out}r+1}{2})}.$
For the bound of $\beta$, please see lemma A.7. ∎
@@
By Gautschi’s inequality (Gautschi, 1959) with $s=\tfrac{1}{2}$ applied at $u-\tfrac{1}{2}$ (so $u>\tfrac{1}{2}$):
@@
Wendel’s inequality (Wendel, 1948) states that for $u>0$ and $s\in(0,1)$,
$\frac{\Gamma(u+s)}{u^{s}\,\Gamma(u)}\;\leq\;1.$